# Eigenstate Thermalization Hypothesis correlations via non-linear Hydrodynamics

### v1

@@ Abstract
Conclusion and Outlook - While free cumulants naturally arise in ETH, it is the classical cumulants that have a hierarchical structure in a hydrodynamic regime. This implies that free cumulants factorize, into products of two-point functions (or a product of two-point functions and one three point function for odd moments, see Eq. (14)). When ETH applies in its general form (1), this provides a simple prediction for the statistics of matrix elements of local operators. These predictions were confirmed in a chaotic one-dimensional spin chain, the spin-1 tilted field Ising model: correlators after the local equilibration time exhibit various power-law decays, with exponent depending on the operator and degree of the correlator. This behavior transitions to an exponential decay at late times, consistent with hydrodynamic predictions.
@@
One aspect of the ETH functions that is not captured by current theories of hydrodynamics is out-of-time ordered correlators of $n\geq 4$ operators, whose Fourier transforms also contribute to Eq. (1). Capturing these would require extending hydrodynamic EFTs beyond their current scope. Another interesting direction to extend our results would be to study higher cumulants in systems exhibiting anomalous transport such as super- or subdiffusion, and in integrable systems with the help of generalized hydrodynamics.
Acknowledgements.
@@
We thank A. Rosch and A. Läuchli for helpful discussions. We also wish to A. Silva for initial discussions on the topic. SP acknowledges funding by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Projektnummer 277101999 - TRR 183 (project B02 and A03), and under Germany’s Excellence Strategy - Cluster of Excellence Matter and Light for Quantum Computing (ML4Q) EXC 2004/1 - 390534769. JW acknowledges support from DFG, under Grant No. 531128043, as well as under Grant No. 397107022, No. 397067869, and No. 397082825 within the DFG Research Unit FOR 2692, under Grant No. 355031190. RM and LD are supported by an NSF award No. PHY2412710.
@@
We briefly introduce the numerical method used in the main text to calculate free, classical and ETH cumulants, which is based on the dynamical quantum typicality (DQT) [59], and check its accuracy by comparing the results to exact diagonalization (ED) results.
@@
According to Eqs. (3) and (7) in the main text, the free cumulants $\kappa_{n}$ and classical cumulants $r_{n}$ are unambiguous determined by $n$-point correlation function $\langle\hat{A}(t_{n-1})\hat{A}(t_{n-2})\cdots\hat{A}(t_{1})\hat{A}\rangle$. In the following, we introduce the numerical method for computing this $n$-point correlation function.
@@
In our simulations, we choose $N_{p}=M\cdot 3^{18-L}$ ($M\geq 1$), which ensure similar accuracy for different system size $L$. Employing two auxiliary states
@@
Figure S1: Accuracy check: DQT (open) vs ED (solid line) for free and classical cumulants for $L=10$ for (a) $\hat{A}=\hat{h}_{i},\beta=0$;(b)$\hat{A}=\hat{s}^{z}_{i},\beta=0$; (c) $\hat{A}=\hat{s}^{x}_{i},\beta=0$; (d) $\hat{A}=\hat{j}_{i},\beta=0$; (e) $\hat{A}=\hat{j}_{\text{tot}},\beta=0$; (f) $\hat{A}=\hat{h}_{i},\beta=0.2$. The DQT results are obtained from a average over $3^{8}$ different realization of Haar random states.
@@
The time evolution of is calculated by iterating in real-time using Chebyshev polynomial algorithm [76, 77]. With similar methods, $|\psi_{\beta}\rangle$ can be calculated by iterating in imaginary time. In this paper, we calculate the free and classical cumulants up to Hilbert space dimension $D=3^{18}\approx 400000000$, far beyond the limit of ED.
Figure S2: Accuracy check: DQT (open) vs ED (solid line) for ETH cumulants for $L=10$ for $\hat{A}=\hat{h}_{i},\beta=0$. The DQT results are obtained from a average over $3^{8}$ different realization of Haar random states.
ETH cumulants
The calculation of ETH cumulants $\kappa^{\text{ETH}}_{n}$ is significantly more involved than that of free or classical cumulants, due to the fact that they are not fixed by $n$-point correlation functions. It is only feasible through dynamical quantum typicality (DQT) in specific cases, e.g., ${\hat{A}}=\hat{h}_{i}$. In the following, we introduce the method to calculate $\kappa^{\text{ETH}}_{2}$ and $\kappa^{\text{ETH}}_{3}$ of $\hat{h}_{i}$.
The second ETH cumulant reads
@@
$=\langle(\hat{A}(t)-\text{Diag}(\hat{A}))(\hat{A}-\text{Diag}(\hat{A}))\rangle,$ (S11)
@@
Similarly, the third ETH cumulant of $\hat{h}_{i}$ can be written as
@@
To check the accuracy of our DQT results, we compare them with those obtained from exact diagonalization (ED) for a system size of $L=10$. The comparison is shown in Fig.S1 for free and classical cumulants, and in Fig.S2 for ETH cumulants. For all operators considered, the DQT and ED results show nearly perfect agreement.

### v2

@@ Abstract
Conclusion and Outlook - While free cumulants naturally arise in ETH, it is the classical cumulants that have a hierarchical structure in a hydrodynamic regime. This implies that free cumulants factorize, into products of two-point functions (or a product of two-point functions and one three point function for odd moments, see Eq. (12)). When ETH applies in its general form (1), this provides a simple prediction for the statistics of matrix elements of local operators. These predictions are supported by numerics in three paradigmatic chaotic one-dimensional systems: correlators after the local equilibration time exhibit various power-law decays, with exponent depending on the operator and degree of the correlator. This behavior transitions to an exponential decay at late times, consistent with hydrodynamic predictions.
@@
One aspect of the ETH functions that is not captured by current theories of hydrodynamics is out-of-time ordered correlators of $n\geq 4$ operators, whose Fourier transforms also contribute to Eq. (1). Capturing these would require extending hydrodynamic EFTs beyond their current scope. Another interesting direction would be to study higher cumulants in systems exhibiting anomalous transport such as super- or subdiffusion, and in integrable systems [78] with the help of generalized hydrodynamics.
@@
Acknowledgment - We thank A. Rosch and A. Läuchli for helpful discussions. We also wish to thank A. Silva for initial discussions on the topic. SP acknowledges funding by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Projektnummer 277101999 - TRR 183 (project B02 and A03), and under Germany’s Excellence Strategy - Cluster of Excellence Matter and Light for Quantum Computing (ML4Q) EXC 2004/1 - 390534769. JW acknowledges support from DFG, under Grant No. 531128043, as well as under Grant No. 397107022, No. 397067869, and No. 397082825 within the DFG Research Unit FOR 2692, under Grant No. 355031190. RM and LD are supported by an NSF award No. PHY2412710. Additionally, we greatly acknowledge computing time on the HPC3 at the University of Osnabrück, granted by the DFG, under Grant No. 456666331.
Data availability - The data that support the findings of this article are openly available [79].
@@
In the Supplemental Material, we present additional numerical results of:
  1. 1.
    Free/classical cumulants for different observables and temperatures in spin-$1$ Ising model;
  2. 2.
    Second ETH free cumulants for energy density observable in the spin-$1$ Ising model;
  3. 3.
    Free/classical cumulants for energy density observables in the spin-$1/2$ Ising model;
  4. 4.
    Details on the numerical methods based on dynamical quantum typicality.
Hydrodynamic free-cumulants for different observables in spin-$1$ Ising model
Figure S1: Local and total current: Free cumulants $\kappa_{n}(t)$ and classical cumulants $r_{n}(t)$ versus time $t$ for local current $\hat{j}_{i}$ [(a)(b)(c)] and total current $\hat{j}_{\text{tot}}$ [(d)(e)(f)] for $L=10,12,14,16,18$ (from light to dark blue). The dashed line indicates the hydrodynamic prediction.
To evaluate the generality of our analytical prediction, we present results for both free and classical cumulants of various observables, including local current $\hat{j}_{i}$ and global current $\hat{j}_{\text{tot}}=\sum_{i}\hat{j}_{i}$ (Fig. S1); $\hat{s}^{z}_{i}$ (Fig. S2) and $\hat{s}^{x}_{i}$ (Fig. S3). Good agreement with our hydrodynamic prediction is observed for most of the observables considered. Deviations are visible in $\kappa_{4}$ and $r_{4}$ for the total current $\hat{j}_{\text{tot}}$ [Fig. S1(e,f)] and in $\kappa_{4}$ for the local spin component $\hat{s}^{x}_{i}$ (Fig. S3 (c)). For $\hat{s}^{x}_{i}$, the deviation in $\kappa_{4}$ is likely due to finite-size effects, as the figure shows that convergence has not been reached within the system sizes considered (in comparison to Fig. S2 (c)). For $\hat{j}_{\text{tot}}$, the finite-size effects are even more pronounced, as a consequence of its collective nature and rapid decay, which make it significantly more difficult to observe the hydrodynamic prediction—particularly for higher cumulants. Nevertheless, it is notable that the predicted $t^{-4.0}$ scaling is still clearly visible in the two-point correlation function.
Moreover, in addition to Fig. 2 (b) in the main text, we present results for the operator $\hat{A}=\hat{h}_{i}$ at finite temperature $\beta=0.2$ for various system sizes, which suggest a convergence toward the hydrodynamic prediction as the system size increases.
Figure S2: Free cumulants $\kappa_{n}(t)$ and classical cumulants $r_{n}(t)$ versus time $t$ for operator ${\hat{A}}=\hat{s}_{i}^{z}$ and $\beta=0$, for $L=10,12,14,16,18$ (from light to dark blue). Blue, black and red dashed line indicate the hydrodynamic prediction $\propto t^{-m/2}$, where $m=1$ in [(a)], $m=2$ in [(b)-(c)] and $m=3$ in (d).
Figure S3: Similar to Fig. S2, but for operator $\hat{A}=\hat{s}^{x}_{i}$.
Figure S4: Free cumulants $\kappa_{n}(t)$ and classical cumulants $r_{n}(t)$ versus time $t$ for operator ${\hat{A}}=\hat{h}_{i}$ and $\beta=0.2$, for $L=10,12,14,16,18,19$ (from light to dark blue). Blue, black and red dashed line indicate the hydrodynamic prediction $\propto t^{-m/2}$, where $m=1$ in [(a)], $m=2$ in [(b)-(c)] and $m=3$ in (d).
Finite-size exponential decay for the second ETH free cumulants
Analogous to Fig. 3 in the main text, Fig. S5 presents results for the second ETH cumulant, $\kappa^{\text{ETH}}_{2}$, of the energy density operator $\hat{A}=\hat{h}_{i}$. Similar behavior is observed: 1) $\kappa^{\text{ETH}}_{2}$ exhibits a power-law decay at intermediate times, followed by exponential behavior $\sim e^{-\Gamma t}$ at later times; 2) the decay rate $\Gamma$ decreases with increasing system size and temperature.
Figure S5: ETH free cumulants $\kappa^{\text{ETH}}_{2}(t)$ as a function of time $t$ for the operator ${\hat{A}}=\hat{h}_{i}$ for (a, b) $\beta=0$ with system sizes $L=10,12,14,16,18,19$ (from light to dark blue), and (c) $L=19$ for $\beta=0,0.1,0.2,0.3,0.4$ (from light to dark blue). The dashed line represents the power-law scaling $\propto t^{-0.5}$, and the dotted line corresponds to the exponential fit $\propto\exp(-\Gamma t)$. The exponential fitting is performed in the region where $\kappa_{2}^{\mathrm{ETH}}(t)\in[10^{-2.5},10^{-4}]$ for $L>12$, and in $\kappa_{2}^{\mathrm{ETH}}(t)\in[10^{-1.5},10^{-2.5}]$ for $L\leq 12$. The corresponding fitting parameters $\Gamma$ are shown in the insets of panel (b) on a double-logarithmic scale and in panel (c) on a linear scale.
Hydrodynamic free cumulants in spin-$\frac{1}{2}$ Ising model
As a further evaluation, we checked our analytical prediction in a spin-$\frac{1}{2}$ mixed field Ising model, with the Hamiltonian $\hat{H}=\sum_{i}\hat{h}_{i}$, where
$\hat{h}_{i}=J\hat{\sigma}_{i}^{z}\hat{\sigma}_{i+1}^{z}+\frac{1}{2}g_{z}(\hat{\sigma}_{i}^{z}+\hat{\sigma}_{i+1}^{z})+\frac{1}{2}g_{x}(\hat{\sigma}_{i}^{x}+\hat{\sigma}_{i+1}^{x})$ (S1)
denotes the energy density. Here $\hat{\sigma}_{i}^{\mu}$ are the Pauli operators in the $\mu=x,z$ direction and we use periodic boundary conditions $\hat{\sigma}_{L+1}^{\mu}=\hat{\sigma}_{1}^{\mu}$. The parameters are fixed as $g_{x}=1.1$, $g_{z}=0.9$ and $J=1.0$. The free and classical cumulants for the local energy density operator $A=\hat{h}_{i}$ are shown in Fig. S6 at infinite temperature $\beta=0$ and in Fig. S7 at finite temperature $\beta=0.2$. A good agreement with the hydrodynamic predictions (given in Eq. (15) in the main text) is observed, with a tendency toward convergence as the system size increases, further supporting the generality of our results.
Figure S6: Free cumulants $\kappa_{n}(t)$ and classical cumulants $r_{n}(t)$ of the energy density $\hat{A}=\hat{h}_{i}$ versus time $t$ in the spin-$\frac{1}{2}$ Ising model: (a) $\kappa_{2}(t)$; (b) $\kappa_{4}(t)$; (c) $\kappa_{3}(t)$ and (d) $r_{4}(t)=\kappa_{4}(t)-(\kappa_{2}(2t))^{2}$ for L=22,24,26,28,30L=22,24,26,28,30(from light to dark blue), at infinite temperature $\beta=0$. Blue, black and red dashed lines indicate the hydrodynamic predictions from Eq. (15) in the main text.
Figure S7: Similar to Fig. S6, but at finite temperature $\beta=0.2$
Details on Dynamical Quantum Typicality
@@
We briefly introduce the numerical method used in the main text to calculate free, classical and ETH free cumulants, which is based on the dynamical quantum typicality (DQT) [58], and check its accuracy by comparing the results to exact diagonalization (ED) results.
@@
According to Eqs. (3) and (7) in the main text, the free cumulants $\kappa_{n}$ and classical cumulants $r_{n}$ are unambiguously determined by $n$-point correlation function $\langle\hat{A}(t_{n-1})\hat{A}(t_{n-2})\cdots\hat{A}(t_{1})\hat{A}\rangle$. In the following, we introduce the numerical method for computing this $n$-point correlation function.
@@
In our simulations, we choose $N_{p}=M\cdot 3^{L_{\text{max}}-L}$ ($M\geq 1$), which ensures similar accuracy for different system size $L$. Employing two auxiliary states
@@
The time evolution of the state is calculated by iterating in real time using Chebyshev polynomial algorithm [85, 86]. With similar methods, $|\psi_{\beta}\rangle$ can be calculated by iterating in imaginary time. In this paper, we calculate the free and classical cumulants up to Hilbert space dimension $D=3^{18}\approx 400000000$, far beyond the limit of ED.
[…]
Figure S9: Accuracy check: DQT (open) vs ED (solid line) for ETH free cumulants (a) $\kappa^{\text{ETH}}_{2}(t)$ and (b) $\kappa^{\text{ETH}}_{3}(t)$ for $L=10$ for $\hat{A}=\hat{h}_{i},\beta=0$. The DQT results are obtained from an average over $3^{8}$ different realizations of Haar random states.
ETH free cumulants
The calculation of ETH free cumulants $\kappa^{\text{ETH}}_{n}$ is significantly more involved than that of free or classical cumulants, due to the fact that they are not fixed by $n$-point correlation functions. It is only feasible through dynamical quantum typicality (DQT) in specific cases, e.g., ${\hat{A}}=\hat{h}_{i}$. In the following, we introduce the method to calculate $\kappa^{\text{ETH}}_{2}$ and $\kappa^{\text{ETH}}_{3}$ of $\hat{h}_{i}$.
The second ETH free cumulant reads
@@
$),\displaystyle=\langle(\hat{A}(t)-\text{Diag}(\hat{A}))(\hat{A}-\text{Diag}(\hat{A}$$)\rangle$, (S12)
@@
Similarly, the third ETH free cumulant of $\hat{h}_{i}$ can be written as
@@
To check the accuracy of our DQT results, we compare them with those obtained from exact diagonalization (ED) for a system size of $L=10$. The comparison is shown in Fig. S8 for free and classical cumulants, and in Fig. S9 for ETH free cumulants. For all operators considered, the DQT and ED results show nearly perfect agreement.